### v1

@@ 4.3 Ablation Study
As a result, training on the MedShape Pelvic dataset performed better than training on PU1k. This indicates that PUCRN has indeed learned Pelvic’s shape features. This is very important. However, Grad PU’s model on the PU1k dataset performed better than MedShape Pelvic’s test set. These reasons are multifaceted, and Grad PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShape Pelvic obtained from real volume medical images.
Table 3: Pelvic point cloud upsampling results in two different train sets.
Methods 512 input 1024 input 2048 input 4096 input
@@
Grad-PU(MedShapeNet) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
PUCRN(MedShapeNet) 3.104 7.806 3.946 16.635 1.839 5.709 1.274 5.087
@@
4.4 Visualization
In addition to the results of point cloud upsampling, what we need to achieve is the combination of medical image segmentation and point cloud upsampling steps to achieve the effect of SSM. In this visualization section, due to the lack of ground truth, we can only choose to display the reconstruction effect. If we want to make quantitative comparisons, it involves the need for No Reference Mesh Quality Assessment. However, mostly 3D quality assessment focusing on point cloud instead of mesh [32].
@@ 4.4 Visualization
Now we input the Pelvic1k dataset into SAM Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds. The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling. The model surface is reconstructed using the Voxel Grid with Marching Cubes[17]. As shows in Figure 2, compared with the surface obtained by directly segmenting voxels through marching cubes (Pseudo Ground Truth), our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN.
@@ 5 Conclusion
This study introduced a framework integrating point cloud upsampling with statistical shape modeling to improve high-resolution pelvic model reconstruction. By combining SAM-Med3D segmentation with upsampling networks, we transformed sparse voxel data into dense point clouds and reconstructed detailed surfaces using the Marching Cube algorithm. Experiments demonstrated the necessity of upsampling and confirmed the framework’s effectiveness in achieving smoother and more complete reconstructions.
Future work will extend this framework to other anatomical structures and incorporate no reference quality assessment for objective evaluations.
@@
5.0.1 Acknowledgements
This research was supported by the National Natural Science Foundation of China under Grant 61773164.
[…]
The Department of Rehabilitation Science, The Hong Kong Polytechnic University, has not financial support, and there are no potential conflicts of interest to disclose.

### v3

@@ 4.3 Ablation Studies
As a result, training on the MedShapePelvic dataset performed better than training on PU1k. This indicates that PUCRN has indeed learned Pelvic’s shape features. This is very important. However, Grad-PU on the PU1k dataset performed better than MedShapePelvic’s test set. These reasons are multifaceted, and Grad-PU may be better able to learn shape features and knowledge from artificially synthesized PU1k datasets, rather than being adept at learning features from MedShapePelvic obtained from real volume medical images.
Table 3: Pelvic point cloud upsampling results in two different train sets: Network(trainset).
Method 512 input 1024 input 2048 input 4096 input
@@
Grad-PU(MedShapePelvic) 5.643 124.639 3.682 82.969 2.477 49.451 1.690 28.871
PUCRN(MedShapePelvic) 3.104 7.806 3.946 16.635 1.839 5.709 1.274 5.087
@@
4.4 Reconstruction Visualizations
In addition to the results of point cloud upsampling, what we need to achieve is the combination of medical image segmentation and point cloud upsampling steps to achieve the effect of SSM.
The Pelvic1k subdataset 3,4 and 6 have been inputted into SAM-Med3D, and after medical image segmentation, we obtain predicted masks of the pelvic bone, which are then stacked to obtain voxels and converted into point clouds. The current point cloud is sparse, and the segmentation point cloud we obtained is standardized to input 65,536 points into the upsampling model and perform 16×16$\times$16 × upsampling. The model surface is reconstructed using the Voxel Grid with Marching Cubes[20], where voxel size is $1.5$. As shows in Figure 4, compared with the surface of ground truth, and obtained by directly segmenting voxels through marching cubes. Our proposed framework achieves smoother after upsampling in both Grad-PU and PUCRN and more complete shape reconstruction in PUCRN. And there are no obvious contours of layer after reconstructed.
4.5 Reconstruction Evaluations
To evaluate the quality of the mesh surface vis-a-vis ground truth, we used three metrics in Section3.4.2 For Surface Reconstruction.
Table 4: Surface reconstruction results on Pelvic1k subdataset by mesh quality metrics: Area-Length Ratio (ALR), Manifoldness Rate (MR), and Connected Component Discrepancy (CC Diff.).
Subdataset 3 Subdataset 4 Subdataset 6
[…]
Ground Truth 0.692 100.0 0.26 0.721 100.0 1.68 0.742 100.0 0.66
Segmented Result (by SAM-Med3D) 0.676 100.0 1.94 0.699 100.0 4.89 0.735 100.0 1.18
PUSSM (by Grad-PU) 0.811 100.0 2.77 0.817 99.9 9.66 0.819 100.0 7.77
PUSSM (by PUCRN) 0.821 100.0 2.06 0.817 100.0 5.72 0.824 100.0 5.34
As shows in Table 4. Our method outperforms ground truth and segmented result in terms of mesh quality and achieves stable reconstruction accuracy. Specifically, our method has uniform grid partitioning (as shown in ALR) and eliminates floating artifacts (as shown in CC Diff.). This improvement is attributed to the finer points and distribution brought about by point cloud upsampling.
Furthermore, we conducted quantitative comparisons on the reconstruction results in order to evaluate surface reconstruction accuracy and feature preservation. We used metrics (CD, F1, NC, ECD, EF1) in Section3.4.2 For Surface Reconstruction.
To ensure a fair and uniform comparison, we employ Poisson disk sampling to generate 16,384 points from each mesh, including the ground truth, segmented results, and upsampled reconstructions. This sampling strategy maintains uniform point distribution and adequately captures both global structures and fine details, facilitating reliable metric computation.
Table 5: Surface reconstruction quantitative results on on Pelvic1k subdataset with 16,384 points by poisson sampling.
Subdataset 3 Subdataset 4 Subdataset 6
[…]
Segmented Result (by SAM-Med3D) 8.379 0.214 0.691 7.420 0.408 56.272 0.126 0.711 54.559 0.250 13.997 0.214 0.716 10.724 0.357
PUSSM (by Grad-PU) 7.200 0.223 0.671 7.326 0.445 53.297 0.109 0.651 50.104 0.254 12.024 0.223 0.658 8.351 0.453
PUSSM (by PUCRN) 7.111 0.229 0.766 6.295 0.464 51.799 0.118 0.678 49.824 0.266 11.620 0.229 0.686 8.184 0.476
As shows in Table 5. Our method, across all three subdatasets, PUCRN consistently outperforms both Grad-PU and the raw SAM-Med3D segmentations. For example, on Subdataset 3, PUCRN achieves a significantly lower CD compared to Grad-PU, along with superior edge preservation as reflected by a lower ECD and a higher EF1-score. Similar trends are observed on Subdataset 4 and Subdataset 6, where PUCRN achieves the best overall performance in terms of both global surface quality and local edge fidelity. These results demonstrate the effectiveness of PUCRN in enhancing the downstream surface reconstruction quality by providing denser and more accurate point distributions.
@@ 5 Conclusion
In this work, we proposed PUSSM, an implicit statistical shape modeling framework that leverages deep point cloud upsampling to reconstruct high-resolution pelvic structures from sparse medical data. Our method avoids the limitations of classical SSMs by eliminating the need for landmark annotations and explicitly modeling shape variation. Experiments demonstrate improved surface quality and anatomical plausibility across multiple evaluation metrics.
Future work will extend this framework to 3D SAM model[15] by points based prompting to integrate it as Chain-of-Thought (CoT)[12] for Segmentation.